A motor winding has a resistance of 80 ohm at room temperature of 20 deg C before switching on to a 230V supply. After 4 hours run the winding resistance is 100 ohm. Find the temperature rise if the material resistance temperature co-effiecent is 1/234.5 deg C.


1
Expert's answer
2021-03-18T03:36:55-0400

Q171582


A motor winding has a resistance of 80 ohms at room temperature of 20 deg C before switching on to a 230V supply. After 4 hours run the winding resistance is 100 ohm. Find the temperature rise if the material resistance temperature co-efficient is 1/234.5 deg C.


Solution:


We are given the resistance of the material at 20 oC.

We will take this as our reference resistance and temperature.

R ref = 80 ohms. T ref = 20 0C.


The effect of temperature on resistance is given by the relation


"R = R_{ ref} [ 1 + \u03b1 ( T - T _{ref} ) ]"


where R ref is the reference resistance, Tref is the temperature at the R ref.

α is the material resistance temperature coefficient.


R is the resistance of the material at temperature T.


We have,

R ref = 80 ohms, T ref = 20 0C.

α = 1/ 234.5 0C.


We have to find the temperature of the material when the resistance observed is 100 ohms.

So R = 100 ohms.


Substitute all this information in the formula we have


"100 ohms = 80 ohms [ 1 + \frac{1}{234.5^0C } ( T - 20 ^0C) ]"


divide both the side by 80 ohms we have


"\frac{100Ω }{80Ω } = [ 1 + \frac{1}{234.5^0C } ( T - 20 ^0C) ]"


"1.25 = 1 + \frac{1}{234.5^0C } ( T - 20 ^0C )"


subtract 1 from both the side we have


"1.25 -1 = 1 + \frac{1}{234.5^0C } ( T - 20 ^0C ) - 1"


"0.25 = \frac{1}{234.5^0C } ( T - 20 ^0C )"



multiply both the side by 234.5 0C, we have


"0.25 * 234.5 ^0C = \frac{1}{234.5^0C } ( T - 20 ^0C ) * 234.5 ^0C"


"0.25 * 234.5 ^0C = T - 20 ^0C"



"58.625 ^0C = T - 20 ^0C"


add 20 0C on both the side we have


T = 58.63 0C + 20 0C = 78.625 0C.


which in 2 significant figure is 79 0C.


Hence the final temperature of the motor winding will be 79 0C.


in question, we are asked the temperature rise.


Temperature rise = Final temperature - initial temperature

= 79 0C - 20 0C

= 59 0C.




Hence the temperature rise of the motor winding is 59 0C.